UT 20160427 LUCI1 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson,
Observer AO: GTaylor (LBTO), EPinna, APuglisi (remote from Italy upon request)
Telescope Operator: GBechetti
LBTO Support: TSargent

Plan for the night

- As before: stability FLAO/LUCI and AFC

Summary

- 2/3 of night used for LBTO engineering (pointing model) due to poor seeing

- A bit of AFC quantification, still some work to do

- LUCI1MOS -acq officially released

Details

All times are UT
Files are luci1.20160427.NNNN.fits

Alexander installed a new GEIRS version in the afternoon delivered by Richard Mathar

02:02 Sunset, Weather: haze, a few scattered cirrus around, T = -3 deg, Hum = 60%, Wind ~ 10m/s...there is hope

02:20 Geno checks pointing and collimates, 1.0" on the guider

Stability PSF

02:45 Preset to Oj287a, OK, RunAO, RSmag=15.6 (?), Bin=4, 66modes, 100Hz, failed (no idea why), resent preset, DIMM started at 1.7", shell RIPed, seeing is too bad, stopped with AO-stuff

MOS-acq LUCI1 - verification

03:00 Preset to M67_MOSacq, OK, sky, src #0002/3, slitview #0004, need to adjust FM4, reso slitview #0005, OK,
resent preset again, sky, src, #0006/7, slitview #0008, aligned, check-ima #0009, nice, #0010 300sec spectrum, DIMM 1.1",
LUCI1 MOS-acq has never been checked by the LUCI-team,
Officially verified that MOS-acq also works with LUCI1

Stability PSF

03:54 Seeing seems to have stabilized, preset to Oj287a again, OK, RunAO, seeing went up again (DIMM from 1" to 2.5"), no chance

AFC

03:39 Preset to NGC4147_AFC, OK; RunAO, RS-mag bin=3, 66 modes, #0011, gnagna, wrong mask, corrected
resent preset, OK, RunAO, RSmag = 13.9, Bin=3, 66 modes, 200Hz, K: 20sec, #0013, GUI hang, xkill
#0014 AFC-acq again, way off, aligned, check 20sec #0015, realigned, check #0016, 300sec SUR #0017, #0018 stopped, no AFC-correction, #0019/20/21 test-images, resent preset, OK, RunAO,
failed to acquire ref (seeing fluke), resent preset once more, OK, RunAO, failed, DIMM 1.1 up to 3.2"
Gave up for the moment being, terrible seeing

04:30 Handover to Dave for pointing model data collection

9:55 Seeing has improved considerably and seems to be stable, switch back to AO

Quantifying AFC

10:00 Preset to M92_AFC, OK, RunAO, RS=10.9, Bin=1, 100Hz, 400modes, DIMM 0.9-1", w/o NCPA,
Doing in each filter 2 times SUR 300sec with 16 samples and 15 times 20sec LIR alternating, inspect for changes of IQ
K: #0048 AFC-acq., , #0049 verif.,OK, #0050 (300sec, SUR), #0051-0065 (15*20sec, LIR), mirror aligned, #0066 (300sec, SUR),
#0067-0081 (15*20sec, LIR)
H: AFC-acq. #0082, #0083/0084 (300sec, SUR, erroneously two images), #0085-0099 (15*20sec, LIR),
#0100 AFC-acq., #0101 (SUR, 300sec, AFC doesn't work, aborted), #0102 AFC-acq., #0103 (300sec, SUR, again, doesn't work),
#0104 AFC-acq. once more, #0105 AFC-verif., #0106 (300sec, SUR) lost AFC during the last 5 samples (see ima),
#0107 AFC-acq., #0108 AFC-veri., #0109 SUR AFC doesn't work (exceed 5pixels) but image looks OK, #0110 AFC-acq.,
#0111 AFC-veri., #0112 SUR generic Arbitrator error at offset, aborted, need to preset again, failed,
Geno corrects pointing, preset, OK, RunAO, bayside timeout, preset again, OK, RunAO, OK, ACF-acq. #0113 (K-band), AFC-veri. #0114 (H-band), SUR #0115 wo AFC, system got completely stuck,
doen't sent corrections to mirrors, restart scheduler, didn't help, need to discuss with Alexander, probably need to restart SW
...more work required !!! ...

12:00 Geno is closing up

-- %USERSIG{JochenHeidt - 2016-04-27}%
Topic revision: r7 - 28 Apr 2016, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback